kittybasil:

ran into an issue calculating... (calculus 2 single var)

kittybasil:

\[\int4x(2x^2+3)^7dx\]\\[du=?\]

kittybasil:

latex is wonky over here \[\rightarrow u=2x^2+3\]\[\rightarrow du=?\]

kittybasil:

@Nnesha sorry, I think I've confused myself again :(

Bob:

I forgot my latex.

Bob:

[\\\\fail_latex]////[]] see ?

kittybasil:

hardcore googling time ↑

Bob:

ikr

Bob:

good thing we can still view openstudy posts...

kittybasil:

it poops out a huge brainly ad what you talkin bout

Bob:

but I just close it ._.

kittybasil:

mine has no X out sign though :c

Bob:

rip

Bob:

click out of the box ?

Bob:

it will prob disappear ?

Bob:

works for me

kittybasil:

maybe it's my browser being lame lol school computer. help with math anyone?

Nnesha:

\[\huge\rm u= \color{Red}{2x^2+3}\]\[\large\rm du= derivative~of~2x^2 + derivative ~of~3 \]

kittybasil:

\[(2x^2+3)=2(2x)+0\]this?

Nnesha:

correct so \[\frac{du}{dx}=2(2x)+0\]

Nnesha:

`du/dx` because we are taking the derivative of u with respect to x

kittybasil:

wait, so adding "dx" is due to implicit differentiation or... ?

Nnesha:

i can check your work. so whenever u finish post the work

Nnesha:

dx is because we are taking the derivative with respect to x

kittybasil:

so then \[du=u'dx\]this?

Nnesha:

yes

kittybasil:

okay I got it then :) thank you!

kittybasil:

okay \[dx=\frac{du}{4x}\]\[\int{4x(2x^2+3)^7}\rightarrow\int4x(u)^7\frac{du}{4x}\rightarrow\int{u^7du}\] @Nnesha I think I'm doing something wrong

kittybasil:

*left out a \(dx\) in the first integral sorry

Nnesha:

that's correct

Nnesha:

now apply the power rule for anti derivative

kittybasil:

\[\int u^7du\]\[\frac{u^8}{8}du\]I feel like this is wrong.

kittybasil:

wait no I forgot + C at the end

Nnesha:

yep so u^8/8 +C

kittybasil:

oh so the \(du\) is because the differentiation wrt \(u\)?

Nnesha:

du represent the derivative of u anti derivative is backwards of Derivative so basically you already have the derivative now you're integrating to find the original function in other words u^8/8+C is anti derivative in terms of u replace u with `2x^2+3` to write it in terms of x

kittybasil:

\[\frac{u^8}{8}+C\rightarrow\frac{(2x^2+3)^8}{8}+C\]

kittybasil:

continue to simplify?

kittybasil:

oops I have to go

Nnesha:

no that's it!